OpenStudy (anonymous):

solve the DE using the exact method: y^2dx - x^2dy=0

OpenStudy (anonymous):

how about the solution of the problems u posted earlier?u got it?

OpenStudy (anonymous):

p(x,y)=y^2 Q(x,y)=-x^2

OpenStudy (anonymous):

yeh got them

OpenStudy (anonymous):

these arent exact......they are spearable but it says to solve using exact method

OpenStudy (anonymous):

but they can be made exact if we multiply the eq with integarting factor

OpenStudy (anonymous):

again the given eq is not exact :)

OpenStudy (anonymous):

yeh but u cant get the integrating factor to be of one variable so that method wont work

OpenStudy (anonymous):

yea right

OpenStudy (anonymous):

so its impossible to solve using the exact methods

OpenStudy (anonymous):

ok lets find the integrating factor

OpenStudy (anonymous):

oh we cant find an integrating factor, darn

OpenStudy (anonymous):

so use seperable equations , i get dx/x^2 = dy/y^2